OpenStudy (anonymous):

help : R*S/T, for R=4 3/4,S=2 2/3, and T=-2

OpenStudy (anonymous):

(19/4) = r, 8/3=s/ so 38*-2/3 = -76/3. hope this helps!

OpenStudy (anonymous):

thanxx